# Cross-Lingual Learning vs. Low-Resource Fine-Tuning: A Case Study with Fact-Checking in Turkish

### 01 Mar 2024 v1

@@ 1.   Introduction
Cross-lingual learning has been studied in related problems such as hate speech detection Stappen et al. (2020), rumor detection Lin et al. (2023), abusive language detection Glavaš et al. (2020) and malicious activity detection on social media Haider et al. (2023). For fact-checking Du et al. (2021) proposed a model that jointly encodes COVID-19-related Chinese and English texts. Additionally, Raja et al. (2023) employed joint training of English and Dravidian news articles and also applied zero-shot transfer learning by fine-tuning with English data and testing on Dravidian data.
@@ 2.   Related Work
In recent years, numerous datasets have emerged for fact-checking and they can be categorized based on how claim statements are obtained. Some studies create claim statements by extracting and manipulating content from source documents such as Wikipedia articles can be categorized as artificial claims Thorne et al. (2018); Jiang et al. (2020); Schuster et al. (2021); Aly et al. (2021); Kim et al. (2023). These studies involve human annotators who systematically generate meaningful claims.
@@
There are limited datasets on multilingual fact-checking. FakeCovid Shahi and Nandini (2020a) includes 5182 multilingual news articles related to COVID-19. DANFEVER Nørregaard and Derczynski (2021), a Danish fact-checking dataset, comprises 6407 claims generated systematically following the FEVER Thorne et al. (2018) approach. Similarly, CsFEVER Ullrich et al. (2023) features 3097 claims in Czech using a similar methodology. Additionally, CHEF Hu et al. (2022) contains 10K claims in Chinese. Furthermore, CT-FCC-18 Barrón-Cedeno et al. (2018) contains political fact-checking claims in both English and Arabic, focusing on the 2016 US Election Campaign debates. X-Fact Gupta and Srikumar (2021) comprises 31189 short statements from fact-checking websites across 25 languages. Lastly, Dravidian_Fake Raja et al. (2023) consists of 26K news articles in four Dravidian languages.
@@
To the best of our knowledge, the only other fact-checking dataset that includes Turkish is X-Fact Gupta and Srikumar (2021) which includes claims and evidence documents in 25 languages. Besides the differences in size of the corpus, their Turkish data diverges from ours in a number of ways. Mainly, our focus in the corpus collection is richer monolingual data, rather than a large coverage of languages. The evidence documents in X-fact are through web searches, rather than crawling directly from the fact-checking site. Although there is some overlap in our sources, our data is also more varied in terms of fact-checking sites and topics of the claims. We also include short summaries provided in justifications and additional metadata. The summaries can be valuable for explainability in fact-checking Atanasova et al. (2020a); Kotonya and Toni (2020b); Stammbach and Ash (2020); Brand et al. (2022). In addition, a semi-automated method is applied to eliminate duplicate claims that we crawled from different sources.
@@ Methods.
Similarly, the authors categorized context-based methods as propagation-based Hartmann et al. (2019); Zhou and Zafarani (2019) and source-based Sitaula et al. (2020). Both methods aim to capture social context to uncover the spread of information. While propagation-based models leverage interactions among users on social media by enhancing the interaction network with additional details like spreaders and publishers, source-based approaches rely on the credibility of sources which can be employed to identify bot accounts on social media.
@@ 3.1.   Dataset for Fact-Checking in Turkish (FCTR)
We crawled 6787 claims from the three Turkish fact-checking websites [^4] https://teyit.org/analiz,
@@
that are also listed on the Duke Reporters’ Lab. [^5] https://reporterslab.org/fact-checking/ Dogrulukpayi and Teyit are also members of the International Fact-Checking Network (IFCN) which is a global community of fact-checkers. Our data collection process involved extracting claim statements, the corresponding evidence presented by the editorial teams, summaries providing justifications which are also written by the editors, veracity labels, website URLs and the publication dates of the URLs.
Claims retrieved from Teyit are summarized using the "findings" section, which provides an overview of the evidence statements. Likewise, when it comes to claims sourced from Dogrula, the summary is derived from the final paragraph within the "evidences" section, encapsulating the key findings. In the case of claims obtained from Dogrulukpayi, the dataset includes a dedicated paragraph following the rating section that encapsulates both the claim and the supporting evidence. This paragraph serves as the summary of these claims. Moreover, unique IDs were assigned to each claim in the dataset.
Claims were also marked as multi-modal if they contained keywords such as "video", "photo" and "image" etc. This classification was made because we recognize that claims featuring such terms require verification not only of their textual content but also of any associated visual or video elements. For example, consider the fact-checked claim presented in Figure 6, which includes an image. In this claim, it was stated that the video shared on social media shows the moments when protesters in France set fire to the Alcazar Library in Marseille during the recent protests. The reviewer who gathered supporting information noted that "According to inverse visual search results, the video is not from Marseille; it’s from the Philippines. The building that caught fire is the Manila Central Post Office." As a result, in order to verify such claims every aspect of evidences should be processed. Since our focus in this study is linguistic aspects of fact-checking, we do not make use of claims that require multimodal processing.
@@
Last but not least, since the claims were collected from three distinct sources, we reviewed the claims to identify candidate duplicate claims. To accomplish this, the BERTScore metric Zhang et al. (2019b) was employed that calculates a similarity score by analyzing the contextual embeddings of individual tokens within claim statements. We set the similarity threshold to 0.85 and execute the metric three times in data source pairs. Subsequently, a manual verification process was conducted to confirm whether the outputs from BERTScore indeed corresponded to duplicate claims.
@@
Veracity Labels Sources Counts
@@
Dogrula, Teyit,
Dogrulukpayi
@@
Dogrula, Teyit,
@@
203
mixed Teyit 109
partially false Dogrulukpayi 72
unproven Teyit 37
half true Dogrula 17
mostly false Dogrula 14
mostly true Dogrula 6
@@ 3.2.   Snopes Dataset
Table 2: Veracity label counts in the Snopes dataset777"other" encompasses the following labels: scam, outdated, misattributed, originated as satire, legend, research in progress, fake, recall, unfounded, legit
@@ 4.   Method
In this study, we fine-tuned the Llama-2 Touvron et al. (2023) model for the veracity prediction task. Llama-2 is an open-source, auto-regressive transformer-based language model that was released by the Meta AI team. It has three variants, with parameter sizes of 7 billion, 13 billion, and 70 billion. Our main rationale for utilizing Llama-2 is that it has a very large and almost up-to-date knowledge base. To be more specific, the pre-training data includes information up to September 2022, while the fine-tuning data is up to June 2023. State-of-the-art language models comprise billions of parameters, demanding large GPU memory resources during fine-tuning for downstream tasks. Additionally, the deployment of such models in real-time applications has become increasingly impractical. Therefore, we adopted parameter-efficient fine-tuning and quantization to make the Llama-2 model fit within our GPU memory constraints without sacrificing information. First, LoRA Hu et al. (2021) introduces a small number of additional parameters and updates their weights while keeping the original parameters frozen. Similarly, QLora Dettmers et al. (2023) employs quantization to the frozen parameters to increase memory efficiency without a significant tradeoff.
@@
Zero-shot prompting is a method of instructing a language model to generate predictions based on a provided prompt template, without the need for specific examples. During this decision-making process, language models can utilize both the knowledge that they acquired during pre-training and the template prompt. Zero-shot prompting proves particularly useful when you have fine-tuned a language model for a related task but lack labeled data for the specific task at hand. On the other hand, providing one or more examples from the intended task as prompts is referred to as few-shot prompting. By presenting these samples within the prompt, the model gains a better understanding of the desired output and its structure. Therefore, it often leads to superior performance compared to zero-shot prompting.
@@ 5.1.   Setup
pt### Instruction: Is the following statement "true" or "false"?### Input:A series of photographs show the skeletal remains of the biblical giant Goliath.### Response:false
@@
The SVM model Cortes and Vapnik (1995) and the multilingual BERT (mBERT) model Devlin et al. (2019) were both trained on the same datasets with identical train-dev-test partitions as a baseline. For the SVM model, we used sparse word and n-gram features weighted by tf-idf. The training instances are weighted with inverse class frequency to counteract the class imbalance, particularly in the case of FCTR100 trials. Similarly, we modified the cross-entropy loss function for the mBERT model. This adaptation took into account the inverse class ratios, causing the models to assign a larger penalty to the majority class compared to the minority class.
Prompt engineering played a critical role in the experiments. Various prompt formats were evaluated and the best results were achieved using the Alpaca prompt template Taori et al. (2023), which is provided in Figure 3. The Llama-2 implementations in the Huggingface’s transformers library  [^8] https://huggingface.co/meta-llama were utilized language models in our transfer learning experiments. Although the Llama-2 language model was primarily pre-trained on English data, we confirmed its proficiency in Turkish as well. Since it was pre-trained on relatively recent data, we preferred Llama-2 in our experiments.
@@ 5.3.   Results
claim 10-fold SVM-snopes 0.651 0.709
claim SVM-snopes 0.695 0.763
claim mBERT-snopes 0.705 0.802
claim Llama-7B-snopes 0.766 0.838
claim Llama-13B-snopes 0.814 0.866
claim Llama-70B-snopes 0.826 0.890
@@
First of all, we conducted fine-tuning of the Llama and baseline models using the Snopes dataset. In all trials, input consisted solely of claim statements, without the inclusion of any supporting evidence. The results are summarized in Table 3. According to the results, the Llama-2 model with 70 billion parameters exhibited the best performance compared to other models. Since no supporting evidence was provided, the models were expected to rely on stylistic features for their predictions. It is noteworthy that the SVM models learned purely from stylistic features. Nevertheless, a substantial performance gap exists between the SVM-snopes and the Llama-2 models. This margin could be attributed to the pre-trained knowledge embedded in Llama-2 models. Moreover, the larger Llama-2 models outperformed Llama-7B, suggesting that Llama-13B and Llama-70B leverage their knowledge better than their smaller variant.
@@
Furthermore, when both the claim statement and the summary (which summarizes the evidence provided by crowd workers) were given as input, the Llama-13B model reached a superior 0.89 and 0.828 F1-macro scores on FCTR500 and FCTR1000 datasets respectively and 0.923 and 0.947 F1-binary scores respectively. These scores were significantly higher compared to training the model with claims alone. The reason why we incorporated summaries as input was to examine whether this additional information improves the models’ capabilities. Notably, the Llama models have limited proficiency in Turkish and we observed poor performance when solely presented with claim statements.
@@
claim 10-fold SVM-fctr500 0.682 0.610
claim SVM-fctr500 0.714 0.709
claim mBERT-fctr500 0.653 0.750
claim Llama-7B-fctr500 0.632 0.765
claim Llama-13B-fctr500 0.635 0.679
claim Llama-70B-fctr500 0.649 0.783
+summary mBERT-fctr500 0.752 0.861
+summary Llama-13B-fctr500 0.890 0.923
@@
claim SVM-fctr1000 0.671 0.842
claim mBERT-fctr1000 0.518 0.797
claim Llama-7B-fctr1000 0.561 0.864
claim Llama-13B-fctr1000 0.642 0.839
+summary mBERT-fctr1000 0.729 0.902
+summary Llama-13B-fctr1000 0.828 0.947
@@
Table 6: Impact of number of inputs
@@
In this experiment, we examined the influence of varying training data quantities on model performance. We maintained consistency by utilizing the identical test set employed in the previous experiment given in Table 4. Table 6 illustrates the consequences of manipulating the quantity of training data when employing the Llama-7B model. According to the results, as the number of training instances increases, the F1-macro score exhibits gradual improvement. However, when we employed 300 and 400 training instances, the model’s performance remained almost constant, with both cases yielding remarkably similar results with only a single instance having a label change in the negative direction. This observation suggests that beyond a certain threshold, additional training instances may not provide significant performance gains, highlighting the presence of a saturation point in the learning curve.
@@ 5.4.   Cross-Lingual Transfer Learning
Neural machine translation is an approach that employs deep learning models to translate a text from a source language to a target language Ranathunga et al. (2023). The transformer-based generative large language models are pre-trained massively in English. Therefore, their performance in other languages may not be equally impressive. To tackle this challenge, we conducted translations of the Turkish fact-checking dataset into English utilizing the ChatGPT API. Table 9 presents the veracity detection results on the translated data. Note that we employed the models fine-tuned on the Snopes dataset.
The results suggest that employing translated claims led to higher success rates for Llama models compared to the few-shot prompting approach. However, the success rate of mBERT was not positively influenced by translation. This phenomenon may be attributed to the differences in pre-training data between Llama models and mBERT. To be more specific, the Llama models were massively trained on English corpora, while the pre-trained data for mBERT might exhibit a more uniform language distribution.

### 22 Mar 2024 v2

@@ 1.   Introduction
Cross-lingual learning has been studied in related problems such as hate speech detection Stappen et al. (2020), rumor detection Lin et al. (2023), abusive language detection Glavaš et al. (2020) and malicious activity detection on social media Haider et al. (2023). For fact-checking, Du et al. (2021) proposed a model that jointly encodes COVID-19-related Chinese and English texts. Additionally, Raja et al. (2023) employed joint training of English and Dravidian news articles and also applied zero-shot transfer learning by fine-tuning with English data and testing on Dravidian data.
@@ 2.   Related Work
In recent years, numerous datasets have emerged for fact-checking and they can be categorized based on how claim statements are obtained. Some studies that create claim statements by extracting and manipulating content from source documents such as Wikipedia articles can be categorized as artificial claims Thorne et al. (2018); Jiang et al. (2020); Schuster et al. (2021); Aly et al. (2021); Kim et al. (2023). These studies involve human annotators who systematically generate meaningful claims.
@@
Fact-checking datasets in languages other than English, and multilingual datasets are limited in comparison to English. FakeCovid Shahi and Nandini (2020) includes 5182 multilingual news articles related to COVID-19. DANFEVER Nørregaard and Derczynski (2021), a Danish fact-checking dataset, comprises 6407 claims generated systematically following the FEVER Thorne et al. (2018) approach. Similarly, CsFEVER Ullrich et al. (2023) features 3097 claims in Czech using a similar methodology. Additionally, CHEF Hu et al. (2022) contains 10K claims in Chinese. Furthermore, CT-FCC-18 Barrón-Cedeno et al. (2018) contains political fact-checking claims in both English and Arabic, focusing on the 2016 US Election Campaign debates. X-Fact Gupta and Srikumar (2021) comprises 31189 short statements from fact-checking websites across 25 languages. Lastly, Dravidian_Fake Raja et al. (2023) consists of 26K news articles in four Dravidian languages.
@@
To the best of our knowledge, the only other fact-checking dataset that includes Turkish is X-Fact Gupta and Srikumar (2021) which includes claims and evidence documents in 25 languages. Besides the differences in size of the corpus, their Turkish data diverges from ours in a number of ways. Mainly, our focus in the corpus collection is richer monolingual data, rather than a large coverage of languages. The evidence documents in X-fact are through web searches, rather than crawling directly from the fact-checking site. Although there is some overlap in our sources, our data is also more varied in terms of fact-checking sites and topics of the claims. We also include short summaries provided in justifications and additional metadata. The summaries can be valuable for explainability in fact-checking Atanasova et al. (2020a); Kotonya and Toni (2020b); Stammbach and Ash (2020); Brand et al. (2022); Cekinel and Karagoz (2024). In addition, a semi-automated method is applied to eliminate duplicate claims that we crawled from different sources.
@@ Methods.
Similarly, the authors categorized context-based methods as propagation-based Hartmann et al. (2019); Zhou and Zafarani (2019) and source-based Sitaula et al. (2020). Both methods aim to capture social context to uncover the spread of information. While propagation-based models leverage interactions among users on social media by enhancing the interaction network with additional details like spreaders and publishers, source-based approaches rely on the credibility of sources which can also be employed to identify bot accounts on social media.
@@ 3.1.   Dataset for Fact-Checking in Turkish (FCTR)
We crawled 6787 claims from the three Turkish fact-checking websites: Teyit, Dogrulukpayi and Dogrula. [^4] https://teyit.org/analiz,
@@
All are listed as fact-checking organizations on the Duke Reporters’ Lab. [^5] https://reporterslab.org/fact-checking/ Dogrulukpayi and Teyit are also members of the International Fact-Checking Network (IFCN) which is a global community of fact-checkers. Our data collection process involved extracting claim statements, the corresponding evidence presented by the editorial teams, summaries providing justifications which are also written by the editors, veracity labels, website URLs and the publication dates of the URLs.
Claims retrieved from Teyit are summarized using the ‘findings’ section, which provides an overview of the evidence statements. Likewise, when it comes to claims sourced from Dogrula, the summary is derived from the final paragraph within the ‘evidences’ section, encapsulating the key findings. In the case of claims obtained from Dogrulukpayi, the dataset includes a dedicated paragraph following the rating section that encapsulates both the claim and the supporting evidence. This paragraph serves as the summary of these claims. Moreover, unique IDs were assigned to each claim in the dataset.
Claims were also marked as multi-modal if they contained keywords such as ‘video’, ‘photo’ and ‘image’ etc. This classification was made because we recognize that claims featuring such terms require verification not only of their textual content but also of any associated visual or video elements. For example, consider the fact-checked claim presented in Figure 6, which includes an image. In this claim, it was stated that the video shared on social media shows the moments when protesters in France set fire to the Alcazar Library in Marseille during the recent protests. The reviewer who gathered supporting information noted that ‘According to inverse visual search results, the video is not from Marseille; it’s from the Philippines. The building that caught fire is the Manila Central Post Office.’ As a result, in order to verify such claims every aspect of evidences should be processed. Since our focus in this study is linguistic aspects of fact-checking, we do not make use of claims that require multimodal processing.
@@
Last but not least, since the claims were collected from three distinct sources, we reviewed the claims to identify candidate duplicate claims. To accomplish this, the BERTScore metric Zhang et al. (2019) was employed that calculates a similarity score by analyzing the contextual embeddings of individual tokens within claim statements. We set the similarity threshold to 0.85 and execute the metric three times in data source pairs. Subsequently, a manual verification process was conducted to confirm whether the outputs from BERTScore indeed corresponded to duplicate claims.
@@
Label
Sources
Counts
@@
Dogrula, Teyit, Dogrulukpayi
@@
Dogrula, Teyit, Dogrulukpayi
203
mixed
Teyit
109
partially false
@@
72
unproven
Teyit
37
half true
Dogrula
17
mostly false
Dogrula
14
mostly true
Dogrula
6
@@ 3.2.   Snopes Dataset
Table 2: Veracity label counts in the Snopes dataset777‘other’ encompasses the following labels: scam, outdated, misattributed, originated as satire, legend, research in progress, fake, recall, unfounded, legit
@@ 4.   Method
In this study, we fine-tuned the LLaMA-2 Touvron et al. (2023) model for the veracity prediction task. Llama-2 is an open-source, auto-regressive transformer-based language model that was released by the Meta AI team. It has three variants, with parameter sizes of 7 billion, 13 billion, and 70 billion. Our main rationale for utilizing Llama-2 is that it has a very large and almost up-to-date knowledge base. To be more specific, the pretraining data includes information up to September 2022, while the fine-tuning data is up to June 2023. State-of-the-art language models comprise billions of parameters, demanding large GPU memory resources during fine-tuning for downstream tasks. Additionally, the deployment of such models in real-time applications has become increasingly impractical. Therefore, we adopted parameter-efficient fine-tuning and quantization to make the Llama-2 model fit within our GPU memory constraints without sacrificing information. First, LoRA Hu et al. (2021) introduces a small number of additional parameters and updates their weights while keeping the original parameters frozen. Similarly, QLora Dettmers et al. (2023) employs quantization to the frozen parameters to increase memory efficiency without a significant trade-off.
@@
Zero-shot prompting is a method of instructing a language model to generate predictions based on a provided prompt template, without the need for specific examples. During this decision-making process, language models can utilize both the knowledge that they acquired during pretraining and the template prompt. Zero-shot prompting proves particularly useful when you have fine-tuned a language model for a related task but lack labeled data for the specific task at hand. On the other hand, providing one or more examples from the intended task as prompts is referred to as few-shot prompting. By presenting these samples within the prompt, the model gains a better understanding of the desired output and its structure. Therefore, it often leads to superior performance compared to zero-shot prompting.
@@ 5.1.   Setup
The SVM model Cortes and Vapnik (1995) and the multilingual BERT (mBERT) model Devlin et al. (2019) were both trained on the same datasets with identical train-dev-test partitions as a baseline. For the SVM model, we used sparse word and n-gram features weighted by tf-idf. The training instances are weighted with inverse class frequency to counteract the class imbalance, particularly in the case of FCTR100 trials. Similarly, we modified the cross-entropy loss function for the mBERT model. This adaptation took into account the inverse class ratios, causing the models to assign a higher penalty to the errors on the minority class compared to the majority class.
Prompt engineering played a critical role in the experiments. Various prompt formats were evaluated and the best results were achieved using the Alpaca prompt template Taori et al. (2023), which is provided in Figure 3. The LLaMA-2 implementations in the Huggingface’s transformers library [^8] https://huggingface.co/meta-llama were utilized language models in our transfer learning experiments. Although the LLaMA-2 language model was primarily pretrained on English data, we confirmed its proficiency in Turkish as well. Since it was pretrained on relatively recent data, we preferred LLaMA-2 in our experiments.
@@ 5.3.   Results
claim 10-fold SVM 0.651 0.709
claim SVM 0.695 0.763
claim mBERT 0.705 0.802
claim LLaMA-7B 0.766 0.838
claim LLaMA-13B 0.814 0.866
claim LLaMA-70B 0.826 0.890
@@
First of all, we conducted fine-tuning of the LLaMA and baseline models using the Snopes dataset. In all trials, input consisted solely of claim statements, without the inclusion of any supporting evidence. The results are summarized in Table 3. According to the results, the LLaMA-2 model with 70 billion parameters exhibited the best performance compared to other models. Since no supporting evidence was provided, the models were expected to rely on stylistic features for their predictions. It is noteworthy that the SVM models learned purely from stylistic features. Nevertheless, a substantial performance gap exists between the SVM and the LLaMA-2 models. This margin could be attributed to the pretrained knowledge embedded in LLaMA-2 models. Moreover, the larger LLaMA-2 models outperformed LLaMA-7B, suggesting that LLaMA-13B and LLaMA-70B leverage their knowledge better than their smaller variant.
@@
Furthermore, when both the claim statement and the summary (which summarizes the evidence provided by crowd workers) were given as input, the LLaMA-13B model reached a superior 0.89 and 0.828 F1-macro scores on FCTR500 and FCTR1000 datasets respectively and 0.923 and 0.947 F1-binary scores respectively. These scores were substantially higher compared to training the model with claims alone. The reason why we incorporated summaries as input was to examine whether this additional information improves the models’ capabilities. Notably, the LLaMA models have limited proficiency in Turkish and we observed poor performance when solely presented with claim statements.
@@
claim 10-fold SVM 0.682 0.610
claim SVM 0.714 0.709
claim mBERT 0.653 0.750
claim LLaMA-7B 0.632 0.765
claim LLaMA-13B 0.635 0.679
claim LLaMA-70B 0.649 0.783
+summary mBERT 0.752 0.861
+summary LLaMA-13B 0.890 0.923
@@
claim SVM 0.671 0.842
claim mBERT 0.518 0.797
claim LLaMA-7B 0.561 0.864
claim LLaMA-13B 0.642 0.839
+summary mBERT 0.729 0.902
+summary LLaMA-13B 0.828 0.947
@@
Table 6: Impact of number of inputs on the FCTR500 data
@@
In this experiment, we examined the influence of varying training data quantities on model performance. We maintained consistency by utilizing the identical test set employed in the previous experiment given in Table 4. Table 6 illustrates the consequences of manipulating the quantity of training data when employing the LLaMA-7B model. According to the results, as the number of training instances increases, the F1-macro score exhibits gradual improvement. However, when we employed 300 and 400 training instances, the model’s performance remained almost constant, with both cases yielding remarkably similar results with only a single instance having a label change in the negative direction. This observation suggests that beyond a certain threshold, additional training instances may not provide substantial performance gains, highlighting the presence of a saturation point in the learning curve.
@@ 5.4.   Cross-Lingual Transfer Learning
Neural machine translation is an approach that employs deep learning models to translate a text from a source language to a target language Ranathunga et al. (2023). The transformer-based generative large language models are pretrained massively in English. Therefore, their performance in other languages may not be equally impressive. To tackle this challenge, we conducted translations of the Turkish fact-checking dataset into English utilizing the ChatGPT API. Table 9 presents the veracity detection results on the translated data. Note that we employed the models fine-tuned on the Snopes dataset.
The results suggest that employing translated claims led to higher success rates for LLaMA models compared to the few-shot prompting approach. However, the success rate of mBERT was not positively influenced by translation. This phenomenon may be attributed to the differences in pretraining data between LLaMA models and mBERT. To be more specific, the LLaMA models were massively trained on English corpora, while the pretrained data for mBERT might exhibit a more uniform language distribution.